OpenStudy (anonymous):

a thermometor reads 93 F .what is the equivalent celcius temperature ? Round your answer to the nearest degree Remember C =5/9( f-32) a.109 b.34 ? c.33 d.61

OpenStudy (anonymous):

If C = (5/9)(f-32) Then just substitute in 93 for 'f' and evaluate the function.

OpenStudy (anonymous):

I.e. C = (5/9)(93-32)

OpenStudy (anonymous):

C=5/9(93-32) =5/9(61) 305/9 34

OpenStudy (anonymous):

yeah